alibaby:

Will pick best response A(−4, 2), B(3, 2), C(3, −5), and D(−4, −2) Calculate the area of the polygon. 28 units squared 38.5 units squared 40.5 units squared 49 units squared

alibaby:

@Logic007

alibaby:

@zepdrix

alibaby:

@psirockin2

alibaby:

@SkyVoltage43 @darkheroine

SkyVoltage43:

is this the full question?

alibaby:

yes

alibaby:

@SkyVoltage43?

Ultrilliam:

is there a image?

SkyVoltage43:

i got this ultri just gimme a bit

alibaby:

no image

alibaby:

@Ultrilliam

SkyVoltage43:

without the image it will be hard but i can do this okay

alibaby:

there is no image and okay i really need this quickk

SkyVoltage43:

5 minutes tops mkay

alibaby:

ok

SkyVoltage43:

If we let the width 3--4=7 be the height of a trapezoid, the the bases will be the heights, delta y, on each side... b1=2--2=4, b2=2--5=7 Since the area of a trapezoid is: A=h(b1+b2)/2 we have: A=7(4+7)/2 A=38.5 u^2

SkyVoltage43:

lol with one minute to spare

alibaby:

XD

KjSaif:

@alibaby u seem familiar :)

KjSaif:

from Os

Rachelp:

I need help

alibaby:

i am @kJSaif

alibaby:

??

Logic007:

Step 1: Plot and label the plots on a graph https://www.desmos.com/calculator/nh8afzbtv6

Logic007:

Step 2: Connect those four points. http://prntscr.com/f6utl5 It looks like a trapezium when u connect those points.

Logic007:

Step 3 : Find the area of trapezium Formula for area of trapezium \(A=\frac{1}{2}(a+b)h\)

Logic007:



Logic007:

Find the distance of AB,AD, and BC.

Logic007:

\(AB=7\) \(AD=4\) \(BC=7\)

Logic007:

\(A=\frac{1}{2}(AD+BC) AB\) Plug the values into the formula

Logic007:

\(A=\frac{1}{2}(4+7)(7)\)

Logic007:

\(A=38.5~units^2\)

Logic007:

@alibaby